OpenStudy (anonymous):

solve the system of equation using substitution y=2x+3 y=3x+1

OpenStudy (tkhunny):

2x+3 = 3x+1 You're almost done.

OpenStudy (anonymous):

Take either one of the equations, y=2x+3 for example, and put the 2x on the other side. Then you have y-2x=3. Now you have to substitute what you know into this equation. You know that y=3x+1 from the other equation. Your new equation should look something like 3x+1-2x=3. You can now solve fro the x and then plug the x value you get in for the x in either one of the original equations and solve for y.